# Crepant Transformation correspondence For Toric Stack Bundles

### v1

@@ Abstract.
We prove a crepant transformation correspondence in genus zero Gromov-Witten theory for toric stack bundles related by crepant wall-crossings of the toric fibers. Specifically, we construct a symplectic transformation that identifies $I$-functions toric stack bundles suitably analytically continued using Mellin-Barnes integral approach. We compare our symplectic transformation with a Fourier-Mukai isomorphism between the $K$-groups.
@@ 0.1. Result
More precisely, let $X_{\pm}$ be toric Deligne-Mumford stacks with $\varphi:X_{+}\dashrightarrow X_{-}$ the birational map arising as the stability parameter crosses a wall. Under certain assumption (see Section 2.3), there is a common toric blow-up $\tilde{X}$ with projective birational morphisms $f_{\pm}:\tilde{X}\rightarrow X_{\pm}$ such that $f_{-}=\varphi\circ f_{+}$ and $f^{\star}_{+}K_{X_{+}}=f^{\star}_{-}K_{X_{-}}$, i.e. $X_{+}$ and $X_{-}$ are $K$-equivalent. In [8], Coates, Iritani and Jiang considered this situation and proved a crepant transformation correspondence for genus 00 descendant Gromov-Witten theory of $\mathcal{X}_{\pm}$. Moreover, they relate this correspondence in Gromov-Witten theory with Fourier-Mukai transform between equivariant derived categories of $X_{\pm}$ studied in [9].
This paper establishes a correspondence in genus 00 Gromov-Witten theory for crepant birational maps of toric stack bundles arising from toric wall-crossings. As described in Section 2.4.1 below, we consider toric stack bundles $\mathcal{P}_{\pm}$ with toric fiber $X_{\pm}$ related by a crepant toric wall-crossing. We have toric stack bundles $\tilde{\mathcal{P}}$ with fiber $\tilde{X}$ and $\mathcal{P}_{0}$ with fiber $X_{0}$, where $\tilde{X}$ is the common toric blow-up of $X_{\pm}$ and $X_{0}$ the common blow-down. We have the following diagram with map $\varphi$ induced from the toric wall-crossing on the fiber, $f_{\pm}$ and $g_{\pm}$ induced from the toric blow-ups and blow-downs on the fiber:
@@
Let ~$\tilde{\mathcal{H}}(\mathcal{P}_{\pm})$ be multi-valued version of Givental’s symplectic vector space for $\mathcal{P}_{\pm}$. There exist a degree-preserving $R_{T}(\!(z^{-1})\!)$-linear symplectic transformation $\mathbb{U}:\tilde{\mathcal{H}}(\mathcal{P}_{-})\rightarrow\tilde{\mathcal{H}}(\mathcal{P}_{+})$ such that:
@@
    (0.1) ${K^{0}_{T}(\mathcal{P}_{-})}$${K^{0}_{T}(\mathcal{P}_{+})}$~${\tilde{\mathcal{H}}(\mathcal{P}_{-})}$~${\tilde{\mathcal{H}}(\mathcal{P}_{+}).}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{{\Psi}_{-}}$$\scriptstyle{{\Psi}_{+}}$$\scriptstyle{\mathbb{U}}$
@@
By [14], the $I$-functions $I_{\pm}$ of $\mathcal{P}_{\pm}$, whose definition is recalled in (3.1), lie on the Lagrangian cones of genus 00 Gromov-Witten theory of $\mathcal{P}_{\pm}$. Therefore, Theorem 0.1 can be viewed as a crepant transformation correspondence for genus 00 Gromov-Witten theory of $\mathcal{P}_{\pm}$.
Our approach to Theorem 0.1 is an adaptation of the method used in [8] for the toric case, replacing toric mirror theorem of [7] by its toric stack bundles counterpart in [14]. In fact, [8] is our main motivation for considering Theorem 0.1 and has been the main source of guidance in our study.
The rest of this paper is organized as follows. In Section 1 we briefly introduce Gromov-Witten theory for Deligne-Mumdford stacks and Givental’s symplectic formalism. In Section 2 we recall the definition and basic results for toric stacks and toric stack bundles, and discuss their toric wall-crossings. In Section 3 we prove Theorem 0.1.
@@ 0.2. Acknowledgment
H.-H. T. is supported in part by Simons foundation collaboration grant.
@@ 1.1. Chen-Ruan Orbifold Cohomology
$H^{k}_{\text{CR}}(X)=\bigoplus_{v\in B:k-2\iota_{v}\in 2\mathbb{Z}}H^{k-2\iota_{v}}(X_{v};\mathbb{C})$
@@
$H^{k}_{\text{CR},T}(X)=\bigoplus_{v\in B:k-2\iota_{v}\in 2\mathbb{Z}}H^{k-2\iota_{v}}_{T}(X_{v};\mathbb{C}).$
$H^{\bullet}_{\text{CR},T}(X)$ carries the $T$-equivariant orbifold Poincaré pairing: for $\alpha,\beta\in H^{\bullet}_{\text{CR},T}(X)$ define
@@
using the Atiyah-Bott localization formula, here $S_{T}$ is the localization of $R_{T}=H^{\bullet}_{T}(\text{pt};\mathbb{C})$ with respect to the set of non-zero homogeneous elements.
@@ 1.2. Gromov-Witten Invariant and Quantum Cohomology
Let $X_{g,n,d}$ be the moduli stack of $n$-pointed genus-$g$ degree-$d$ orbifold stable maps to $X$, where $d\in H_{2}(\lvert X\rvert;\mathbb{Z})$. Let $\text{ev}_{i}:X_{g,n,d}\rightarrow IX$ be the evaluation map at the $i$th marked point, and $\psi_{i}\in H^{2}_{T}(X_{g,n,d})$ the equivariant first Chern class of the $i$-th universal cotangent line bundle over $X_{g,n,d}$. When $X$ admits a $T$-action, $X_{g,n,d}$ carries an induced $T$-action. Consider the $T$-equivariant weighted virtual fundamental class $[X_{g,n,d}]^{\text{vir}}\in H_{\bullet,T}(X_{g,n,d};\mathbb{Q})$, see e.g. [1] for details.
@@
We introduce generating functions of Gromov-Witten invariants. Let $\text{NE}(X)$ be the cone in $H_{2}(\lvert X\rvert;\mathbb{R})$ generated by effective curves classes and set
@@ 1.3. Equivariant Quantum Cohomology
Let $\{\phi_{i}\}$ be a basis of $H_{\text{CR}}^{\bullet}(X)$. Let $\eta_{ij}=(\phi_{i},\phi_{j})_{\text{CR}}$ and $\eta^{ij}$ the entry of inverse matrix of $(\eta_{ij})$. For $\tau\in H_{\text{CR}}^{\bullet}(X)$ we define the quantum product at $\tau$ by
@@
For equivariant orbifold cohomology $H^{\bullet}_{\text{CR},T}(X)$ and $\tau\in H^{\bullet}_{\text{CR},T}(X)$ we define the quantum product analogously, so
@@ 1.4. Givental’s Symplectic Formalism
for any $f,g\in\mathcal{H}$, i.e. the coefficient of $z^{-1}$ of the orbifold Poincaré pairing of $f(-z)$ and $g(z)$. This space has a standard polarization
$\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$
@@
$\mathcal{H}_{+}=H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}S_{T}[z]\llbracket Q\rrbracket$
@@
$\mathcal{L}_{X}=\left\{(\mathbf{p},\mathbf{q})\in\mathcal{H}_{+}\oplus\mathcal{H}_{-}:\mathbf{p}=d_{\mathbf{q}}\mathcal{F}^{0}_{X}\right\}\subset T^{*}\mathcal{H}_{+}\cong\mathcal{H}$
@@
$-1z+\mathbf{t}+\sum_{n=0}^{\infty}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{i}\frac{Q^{d}}{n!}\left\langle\mathbf{t},\dots,\mathbf{t},\frac{\phi_{i}}{-z-\psi}\right\rangle^{X}_{0,n+1,d}\phi^{i}$
@@
$J_{X}(\mathbf{t},z)=1z+\mathbf{t}+\sum_{n=0}^{\infty}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{i}\frac{Q^{d}}{n!}\left\langle\mathbf{t},\dots,\mathbf{t},\frac{\phi_{i}}{z-\psi}\right\rangle^{X}_{0,n+1,d}\phi^{i}$
@@
$\nabla_{v}f(\tau)=\partial_{v}f(\tau)+z^{-1}v\bullet_{\tau}f(\tau)$
and denote $\nabla_{\phi_{i}}$ by $\nabla_{i}$. Then the fundamental solution
@@
is determined by following conditions
$\nabla_{i}L(\tau,z)\phi$ $=0$
$\left(vQ\frac{\partial}{\partial Q}-\partial_{v}\right)L(\tau,z)\phi$ $=L(\tau,z)\frac{v}{z}\phi$
[…]
for $i$ runs over basis $\{\phi_{i}\}$ of $H^{\bullet}_{\text{CR},T}(X)$, $\phi\in H^{\bullet}_{\text{CR},T}(X)$ and $v\in H^{2}_{T}(X)$.
@@
$L(\tau,z)\phi_{i}=\phi_{i}+\sum_{j}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{n=0}^{\infty}\frac{Q^{d}}{n!}\left\langle\frac{\phi_{i}}{-z-\psi},\tau,\dots,\tau,\phi_{j}\right\rangle^{X}_{0,n+2,d}\phi^{j}$
@@ 2.1. Toric Variety Associated To A Fan
$H_{m}=\left\{u\in N_{\mathbb{R}}:\langle m,u\rangle=0\right\}$
@@
Geometric information of a toric variety can be deduced from the combinatorial data of its fan. For example, $X_{\Sigma}$ is a smooth if and only if every cone $\sigma\in\Sigma$ is smooth; $X_{\Sigma}$ is an orbifold if and only if every cone $\sigma\in\Sigma$ is simplicial; $X_{\Sigma}$ is compact if and only if $\Sigma$ is complete, i.e. the support $|\Sigma|=\bigcup_{\sigma\in\Sigma}\sigma$ is the entire $N_{\mathbb{R}}$.
@@ 2.1.1. Torus Fixed Points and Toric Blow-ups
Let $\sigma$ be a cone in $\Sigma$, $N_{\sigma}$ the sublattice of $N$ spanned by the points in $\sigma\cap\mathbf{N}$ and $N(\sigma)=N/N_{\sigma}$. The cones $\sigma\in\Sigma$ are in bijective correspondence with $T_{N}$-orbits $O(\sigma)$ of $X_{\Sigma}$. Recall the definition of $U_{\sigma}$, a point $p\in U_{\sigma}$ can be represented by a map $\gamma:S_{\sigma}\rightarrow\mathbb{C}$ which sends $m\in S_{\sigma}$ to $\chi^{m}(p)\in\mathbb{C}$. The torus orbit
$O(\sigma)$ =$=\left\{\gamma:S_{\sigma}\rightarrow\mathbb{C}|\gamma(m)\neq 0\Leftrightarrow m\in\sigma^{\perp}\cap M\right\}\simeq\text{Hom}(\sigma^{\perp}\cap M,\mathbb{C}^{\times})\simeq T_{N(\sigma)}$
In particular, the top-dimensional cone corresponds to torus-fixed points of $X_{\Sigma}$.
@@
$\Sigma^{*}(\sigma)=(\Sigma\setminus\{\sigma\})\cup\Sigma^{\prime}(\sigma)$
@@
Similarly, we can define a toric blow-up along a subvariety $V(\tau)$ using the star subdivision of $\Sigma$ along $\tau$. Let $\tau\in\Sigma$ such that all cones of $\Sigma$ containing $\tau$ are smooth. Let $u_{\tau}=\sum_{\rho\in\tau(1)}u_{\rho}$ and for each $\sigma\in\Sigma$ containing $\tau$, define
$\Sigma^{*}_{\sigma}(\tau)=\left\{\text{Cone}(A):A\subseteq\{u_{\tau}\}\cup\sigma(1),\tau(1)\nsubseteq A\right\}$
@@ 2.2. Toric Stacks
Given $K\cong(\mathbb{C}^{\times})^{r}$, characters of $K$
$D_{1},\dots,D_{m}\in\mathbb{L}^{\vee}=\text{Hom}(K,\mathbb{C}^{\times})$
define a map from $K$ to (×$(\mathbb{C}^{\times})^{m}$, hence an action of $K$ on $\mathbb{C}^{m}$. For a set $I\subset[m]=\{1,2,\dots,m\}$, define
$\angle_{I}$ =$=\left\{\sum_{i\in I}a_{i}D_{i}:a_{i}\in\mathbb{R},a_{i}>0\right\}\subset\mathbb{L}^{\vee}\otimes\mathbb{R},$
(×$(\mathbb{C}^{\times})^{I}\times(\mathbb{C})^{\overline{I}}$ =$=\left\{(z_{1},\dots,z_{m}):z_{i}\neq 0\ \text{for}\ i\in I\right\}\subset\mathbb{C}^{m}$
where $\overline{I}=[m]\setminus I$ is the complement of $I$. Set =$\angle_{\emptyset}=\{0\}$.
@@
with the assumptions that $[m]\in\mathcal{A}_{\omega}$ and for each $I\in\mathcal{A}_{\omega}$, $\{D_{i}:i\in I\}$ spans $\mathbb{L}^{\vee}\otimes\mathbb{R}$ over $\mathbb{R}$.
Let $S\subset[m]$ be the set of indices $i$ such that $[m]\setminus\{i\}\notin\mathcal{A}_{\omega}$. Then the characters $\{D_{i}:i\in S\}$ are linearly independent and every $I\in\mathcal{A}_{\omega}$ contains $S$. In fact we may define $X_{\omega}=[U^{\prime}_{\omega}/G]$ where $\mathcal{A}_{\omega}=\{I\sqcup S:I\in\mathcal{A}^{\prime}_{\omega}\}$, $U_{\omega}\cong U^{\prime}_{\omega}\times(\mathbb{C}^{\times})^{|S|}$, $G=\text{Ker}(K\rightarrow(\mathbb{C}^{\times})^{|S|})$ and get the same $X_{\omega}$. And this is the original construction of toric Deligne-Mumford stacks of Borisov-Chen-Smith [3].
The definition here is actually equivalent to the construction by Jiang using an $S$-extended stacky fan [12]. An $S$-extended stacky fan $\mathbf{\Sigma}=(\mathbf{N},\Sigma,\beta,S)$ consists the following data: a finitely generated abelian group $\mathbf{N}$, a rational simplicial fan $\Sigma$ in $\mathbf{N}\otimes\mathbb{R}$, a homomorphism $\beta:\mathbb{Z}^{m}\rightarrow\mathbf{N}$ with $b_{i}=\beta(e_{i})\in\mathbf{N}$ and $\overline{b_{i}}$ the image of $b_{i}$ in $\mathbf{N}\otimes\mathbb{R}$, and $S\subset[m]$ such that $1$-dimensional cones in $\Sigma$ are in bijective correspondence with $\overline{b_{i}}$ for $i\in[m]\setminus S$, and for each $i\in S$, $\overline{b_{i}}$ lies in the support $\lvert\Sigma\rvert$.
@@
[…]
i.e. define $\beta:\mathbb{Z}^{m}\rightarrow\mathbf{N}$ to be the cokernel of $\mathbb{L}\rightarrow\mathbb{Z}^{m}$. Let $b_{i}=\beta(e_{i})$ and define $S$ as the set of indices $i$ such that $[m]\setminus\{i\}\notin\mathcal{A}_{\omega}$. Then $\Sigma_{\omega}$ is the fan generated by elements $\{\overline{b_{i}},i\in[m]\setminus S\}$ and we get the $S$-extended stacky fan $\mathbf{\Sigma}=(\mathbf{N},\Sigma_{\omega},\beta,S)$.
@@
$0\rightarrow\mathbf{N}^{\vee}\longrightarrow(\mathbb{Z}^{m})^{\vee}\longrightarrow\mathbb{L}^{\vee}$
@@
$\mathcal{A}_{\omega}=\{I\subset[m]:S\subset I\ \text{and}\ \sigma_{\overline{I}}\ \text{is a cone of}\ \Sigma\}$

### v2

@@ Abstract.
We prove a crepant transformation correspondence in genus zero Gromov-Witten theory for toric stack bundles related by crepant wall-crossings of the toric fibers. Specifically, we construct a symplectic transformation that identifies $I$-functions of toric stack bundles suitably analytically continued using Mellin-Barnes type integrals. We compare our symplectic transformation with a Fourier-Mukai isomorphism between the $K$-groups.
@@ 0.1. Result
More precisely, let $X_{\pm}$ be toric Deligne-Mumford stacks with $\varphi:X_{+}\dashrightarrow X_{-}$ the birational map arising as the stability parameter crosses a wall. Under certain assumption (see Section 2.3), there is a common toric blow-up $\tilde{X}$ with projective birational morphisms $f_{\pm}:\tilde{X}\rightarrow X_{\pm}$ such that $f_{-}=\varphi\circ f_{+}$ and $f^{\star}_{+}K_{X_{+}}=f^{\star}_{-}K_{X_{-}}$, i.e. $X_{+}$ and $X_{-}$ are $K$-equivalent. In [8], Coates, Iritani and Jiang considered this situation and proved a crepant transformation correspondence for genus 0 descendant Gromov-Witten theory of $X_{\pm}$. Moreover, they relate this correspondence in Gromov-Witten theory with Fourier-Mukai transform between equivariant derived categories of $X_{\pm}$ studied in [9].
This paper establishes a correspondence in genus 0 Gromov-Witten theory for crepant birational maps of toric stack bundles arising from toric wall-crossings. As described in Section 2.4.1 below, we consider toric stack bundles $\mathcal{P}_{\pm}$ with toric fiber $X_{\pm}$ related by a crepant toric wall-crossing. We have toric stack bundles $\tilde{\mathcal{P}}$ with fiber $\tilde{X}$ and $\mathcal{P}_{0}$ with fiber $X_{0}$, where $\tilde{X}$ is the common toric blow-up of $X_{\pm}$ and $X_{0}$ is the common blow-down. We have the following diagram with map $\varphi$ induced from the toric wall-crossing on the fiber, $f_{\pm}$ and $g_{\pm}$ induced from the toric blow-ups and blow-downs on the fiber:
@@
Let $\tilde{\mathcal{H}}(\mathcal{P}_{\pm})$ be multi-valued version of Givental’s symplectic vector space for $\mathcal{P}_{\pm}$. There exist a degree-preserving $R_{T}(\!(z^{-1})\!)$-linear symplectic transformation $\mathbb{U}:\tilde{\mathcal{H}}(\mathcal{P}_{-})\rightarrow\tilde{\mathcal{H}}(\mathcal{P}_{+})$ such that:
@@
    (0.1) $K^{0}_{T}(\mathcal{P}_{-})$$K^{0}_{T}(\mathcal{P}_{+})$$\tilde{\mathcal{H}}(\mathcal{P}_{-})$$\tilde{\mathcal{H}}(\mathcal{P}_{+}).$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{{\Psi}_{-}}$$\scriptstyle{{\Psi}_{+}}$$\scriptstyle{\mathbb{U}}$
@@
By [14], the $I$-functions $I_{\pm}$ of $\mathcal{P}_{\pm}$, whose definition is recalled in (3.1), lie on the Lagrangian cones of genus 0 Gromov-Witten theory of $\mathcal{P}_{\pm}$. Therefore, Theorem 0.1 can be viewed as a crepant transformation correspondence for genus 0 Gromov-Witten theory of $\mathcal{P}_{\pm}$.
Our approach to Theorem 0.1 is an adaptation of the method used in [8] for the toric case, replacing the toric mirror theorem of [7] by its toric stack bundles counterpart in [14]. In fact, [8] is our main motivation for considering Theorem 0.1 and has been the main source of guidance in our study.
The rest of this paper is organized as follows. In Section 1 we briefly introduce Gromov-Witten theory for Deligne-Mumdford stacks and Givental’s symplectic formalism. In Section 2 we recall the definitions and basic results for toric stacks and toric stack bundles, and discuss their toric wall-crossings. In Section 3 we prove Theorem 0.1.
@@ 0.2. Acknowledgment
We thank the referee for helpful comments and suggestions. H.-H. T. is supported in part by Simons foundation collaboration grant.
@@ 1.1. Chen-Ruan Orbifold Cohomology
$H^{k}_{\text{CR}}(X)=\bigoplus_{v\in\mathsf{B}:k-2\iota_{v}\in 2\mathbb{Z}}H^{k-2\iota_{v}}(X_{v};\mathbb{C})$
@@
$H^{k}_{\text{CR},T}(X)=\bigoplus_{v\in\mathsf{B}:k-2\iota_{v}\in 2\mathbb{Z}}H^{k-2\iota_{v}}_{T}(X_{v};\mathbb{C}).$
$H^{\bullet}_{\text{CR},T}(X)$ carries the $T$-equivariant orbifold Poincaré pairing: for $\alpha,\beta\in H^{\bullet}_{\text{CR},T}(X)$, define
@@
using the Atiyah-Bott localization formula. Here $S_{T}$ is the localization of $R_{T}=H^{\bullet}_{T}(\text{pt};\mathbb{C})$ with respect to the set of non-zero homogeneous elements.
@@ 1.2. Gromov-Witten Invariant and Quantum Cohomology
[…]
@@
[…]
@@ 1.3. Equivariant Quantum Cohomology
Let $\{\phi_{i}\}$ be a basis of $H_{\text{CR}}^{\bullet}(X)$. Let $\eta_{ij}=(\phi_{i},\phi_{j})_{\text{CR}}$ and $\eta^{ij}$ be the entries of inverse matrix of $(\eta_{ij})$. For $\tau\in H_{\text{CR}}^{\bullet}(X)$, we define the quantum product at $\tau$ by
@@
For equivariant orbifold cohomology $H^{\bullet}_{\text{CR},T}(X)$ and $\tau\in H^{\bullet}_{\text{CR},T}(X)$, we define the quantum product analogously, so
@@ 1.4. Givental’s Symplectic Formalism
for any $f,g\in\mathcal{H}$, i.e. the coefficient of $z^{-1}$ of the orbifold Poincaré pairing of f(z)f(-z) and $g(z)$. This space has a standard polarization
$\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-},$
@@
$\mathcal{H}_{+}=H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}S_{T}[z]\llbracket Q\rrbracket,$
@@
$\mathcal{L}_{X}=\left\{(\mathbf{p},\mathbf{q})\in\mathcal{H}_{+}\oplus\mathcal{H}_{-}:\mathbf{p}=d_{\mathbf{q}}\mathcal{F}^{0}_{X}\right\}\subset T^{*}\mathcal{H}_{+}\cong\mathcal{H}.$
@@
$1z+\mathbf{t}+\sum_{n=0}^{\infty}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{i}\frac{Q^{d}}{n!}\left\langle\mathbf{t},\dots,\mathbf{t},\frac{\phi_{i}}{-z-\psi}\right\rangle^{X}_{0,n+1,d}\phi^{i}$
@@
$J_{X}(\mathbf{t},z)=1z+\mathbf{t}+\sum_{n=0}^{\infty}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{i}\frac{Q^{d}}{n!}\left\langle\mathbf{t},\dots,\mathbf{t},\frac{\phi_{i}}{z-\psi}\right\rangle^{X}_{0,n+1,d}\phi^{i}.$
@@
$\nabla_{v}f(\tau)=\partial_{v}f(\tau)+z^{-1}v\bullet_{\tau}f(\tau).$
Denote $\nabla_{\phi_{i}}$ by $\nabla_{i}$. Then the fundamental solution
@@
is determined by the following conditions
$\nabla_{i}L(\tau,z)\phi$ $=0,$
$\left(vQ\frac{\partial}{\partial Q}-\partial_{v}\right)L(\tau,z)\phi$ $=L(\tau,z)\frac{v}{z}\phi,$
$L(\tau,z)|_{\tau=Q=0}$ $=\text{id},$
for $i$ runs over the basis $\{\phi_{i}\}$ of $H^{\bullet}_{\text{CR},T}(X)$, $\phi\in H^{\bullet}_{\text{CR},T}(X)$, and $v\in H^{2}_{T}(X)$.
@@
$L(\tau,z)\phi_{i}=\phi_{i}+\sum_{j}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{n=0}^{\infty}\frac{Q^{d}}{n!}\left\langle\frac{\phi_{i}}{-z-\psi},\tau,\dots,\tau,\phi_{j}\right\rangle^{X}_{0,n+2,d}\phi^{j}.$
@@ 2.1. Toric Variety Associated To A Fan
$H_{m}=\left\{u\in N_{\mathbb{R}}:\langle m,u\rangle=0\right\}.$
@@
Geometric information of a toric variety can be deduced from the combinatorial data of its fan. For example, $X_{\Sigma}$ is smooth if and only if every cone $\sigma\in\Sigma$ is smooth; $X_{\Sigma}$ is an orbifold if and only if every cone $\sigma\in\Sigma$ is simplicial; $X_{\Sigma}$ is compact if and only if $\Sigma$ is complete, i.e. the support $|\Sigma|=\bigcup_{\sigma\in\Sigma}\sigma$ is the entire $N_{\mathbb{R}}$.
@@ 2.1.1. Torus Fixed Points and Toric Blow-ups
Let $\sigma$ be a cone in $\Sigma$, $N_{\sigma}$ the sublattice of $N$ spanned by the points in $\sigma\cap\mathbf{N}$ and $N(\sigma)=N/N_{\sigma}$. The cones $\sigma\in\Sigma$ are in bijective correspondence with $T_{N}$-orbits $O(\sigma)$ of $X_{\Sigma}$. Recall the definition of $U_{\sigma}$, a point $p\in U_{\sigma}$ can be represented by a map $\gamma:S_{\sigma}\rightarrow\mathbb{C}$ which sends $m\in S_{\sigma}$ to $\chi^{m}(p)\in\mathbb{C}$. The torus orbit is
$O(\sigma)$ $=\left\{\gamma:S_{\sigma}\rightarrow\mathbb{C}|\gamma(m)\neq 0\Leftrightarrow m\in\sigma^{\perp}\cap M\right\}\simeq\text{Hom}(\sigma^{\perp}\cap M,\mathbb{C}^{\times})\simeq T_{N(\sigma)}.$
In particular, a top-dimensional cone corresponds to a torus-fixed point of $X_{\Sigma}$.
@@
$\Sigma^{*}(\sigma)=(\Sigma\setminus\{\sigma\})\cup\Sigma^{\prime}(\sigma),$
@@
Similarly, we can define a toric blow-up along a subvariety $V(\tau)$ using the star subdivision of $\Sigma$ along $\tau$. Let $\tau\in\Sigma$ be such that all cones of $\Sigma$ containing $\tau$ are smooth. Let $u_{\tau}=\sum_{\rho\in\tau(1)}u_{\rho}$ and for each $\sigma\in\Sigma$ containing $\tau$, define
$\Sigma^{*}_{\sigma}(\tau)=\left\{\text{Cone}(A):A\subseteq\{u_{\tau}\}\cup\sigma(1),\tau(1)\nsubseteq A\right\},$
@@ 2.2. Toric Stacks
Given $K\cong(\mathbb{C}^{\times})^{r}$, characters of $K$,
$D_{1},\dots,D_{m}\in\mathbb{L}^{\vee}=\text{Hom}(K,\mathbb{C}^{\times}),$
define a map from $K$ to $(\mathbb{C}^{\times})^{m}$, hence an action of $K$ on $\mathbb{C}^{m}$. For a set $I\subset[m]=\{1,2,\dots,m\}$, define
$\angle_{I}$ $=\left\{\sum_{i\in I}a_{i}D_{i}:a_{i}\in\mathbb{R},a_{i}>0\right\}\subset\mathbb{L}^{\vee}\otimes\mathbb{R},$
$(\mathbb{C}^{\times})^{I}\times(\mathbb{C})^{\overline{I}}$ $=\left\{(z_{1},\dots,z_{m}):z_{i}\neq 0\ \text{for}\ i\in I\right\}\subset\mathbb{C}^{m},$
where $\overline{I}=[m]\setminus I$ is the complement of $I$. Set $\angle_{\emptyset}=\{0\}$.
@@
with the assumptions that $m]\in\mathcal{A}_{\omega}$ and for each $I\in\mathcal{A}_{\omega}$, $\{D_{i}:i\in I\}$ spans $\mathbb{L}^{\vee}\otimes\mathbb{R}$ over $\mathbb{R}$.
Let $S\subset[m]$ be the set of indices $i$ such that $m]\setminus\{i\}\notin\mathcal{A}_{\omega}$. Then the characters $\{D_{i}:i\in S\}$ are linearly independent and every $I\in\mathcal{A}_{\omega}$ contains $S$. In fact we may define $X_{\omega}=[U^{\prime}_{\omega}/G]$ where $\mathcal{A}_{\omega}=\{I\sqcup S:I\in\mathcal{A}^{\prime}_{\omega}\}$, $U_{\omega}\cong U^{\prime}_{\omega}\times(\mathbb{C}^{\times})^{|S|}$, $G=\text{Ker}(K\rightarrow(\mathbb{C}^{\times})^{|S|})$ and get the same $X_{\omega}$. And this is the original construction of toric Deligne-Mumford stacks of Borisov-Chen-Smith [3].
The definition here is actually equivalent to the construction by Jiang using an $S$-extended stacky fan [12]. An $S$-extended stacky fan $\mathbf{\Sigma}=(\mathbf{N},\Sigma,\beta,S)$ consists of the following data: a finitely generated abelian group $\mathbf{N}$, a rational simplicial fan $\Sigma$ in $\mathbf{N}\otimes\mathbb{R}$, a homomorphism $\beta:\mathbb{Z}^{m}\rightarrow\mathbf{N}$ with $b_{i}=\beta(e_{i})\in\mathbf{N}$ and $\overline{b_{i}}$ the image of $b_{i}$ in $\mathbf{N}\otimes\mathbb{R}$, and $S\subset[m]$ such that $1$-dimensional cones in $\Sigma$ are in bijective correspondence with $\overline{b_{i}}$ for $i\in[m]\setminus S$, and for each $i\in S$, $\overline{b_{i}}$ lies in the support $\lvert\Sigma\rvert$.
@@
$0\rightarrow\mathbb{L}\overset{(D_{1},\dots,D_{m})}{\xrightarrow{\hskip 28.45274pt}}\mathbb{Z}^{m}\overset{\beta}{\longrightarrow}\mathbf{N}\rightarrow 0,$
i.e. define $\beta:\mathbb{Z}^{m}\rightarrow\mathbf{N}$ to be the cokernel of $\mathbb{L}\rightarrow\mathbb{Z}^{m}$. Let $b_{i}=\beta(e_{i})$ and define $S$ as the set of indices $i$ such that $m]\setminus\{i\}\notin\mathcal{A}_{\omega}$. Then $\Sigma_{\omega}$ is the fan generated by elements $\{\overline{b_{i}},i\in[m]\setminus S\}$ and we get the $S$-extended stacky fan $\mathbf{\Sigma}=(\mathbf{N},\Sigma_{\omega},\beta,S)$.
@@
$0\rightarrow\mathbf{N}^{\vee}\longrightarrow(\mathbb{Z}^{m})^{\vee}\longrightarrow\mathbb{L}^{\vee},$
@@
$\mathcal{A}_{\omega}=\{I\subset[m]:S\subset I\ \text{and}\ \sigma_{\overline{I}}\ \text{is a cone of}\ \Sigma\},$